OpenStudy (anonymous):

A Stock price increased by 20% on Monday, decreased by 10% on Tuesday and increased by 1/6 on Wednesday. What was the percentage increase from Monday to Wednesday? Is it 26%

OpenStudy (anonymous):

increase: 20%-10%+(1/6*100)%=(80/3)%, approximately 26.67%

OpenStudy (rulnick):

It is exactly 26%.

OpenStudy (rulnick):

1(1+0.2)(1-0.1)(1+1/6) = 1(1.2)(0.9)(7/6) = (6/5)(9/10)(7/6) = 63/50 = 126/100 = 1.26, or an increase of 26%.